weBwk 404 No- Chapte | Chapte | € 01.Scie -Conten, I nib8818user nib818key sualYktnQHrlbesi nm pivot.utsa edu/webworkz/MAT,1073,Fall 2018/Module 3.,Homework/6/effectiveUser write a formula for the graphed function by using transformations of the graphs of one of the toolkit functions h(z) g(z) Click on a graph to enlarge a) a f(z) b. g(z) c. h() For addtional help with this problem type, accoss the following resources TEXT Read Example 8 in Section 1.5 of the text NOTES Review the Try It problems under Graphing Functions Using Vertical a VIDEO Watch this video on finding and evaiuating iers fu VINEn Watch this and Horizontal Shifts in your 1.5 Guided Notes evaluating inverse functions 凸 F8 F9 F10 F11 F12 F5 Fe
media%2F685%2F685e831f-c453-4524-82fa-53
media%2Feb0%2Feb0d5661-e041-4005-b150-d8
media%2F0ef%2F0efa5c1c-753e-4c36-be3d-08
media%2F8a6%2F8a678509-1f7a-46bc-9277-ba
media%2Ff9d%2Ff9dcdd61-0e69-4a38-bf8d-b7
media%2Fb0a%2Fb0a3001f-861f-4095-ae38-f5
media%2F874%2F874520ad-b054-4f8a-86af-4a
media%2F2b9%2F2b92a458-9012-4940-9b86-a3
media%2Fc44%2Fc448b9f6-887a-43fa-a119-95
media%2F016%2F01678e9e-fc6a-4198-bd6e-46
media%2F88c%2F88c2416a-00a4-45ea-bb3a-50
media%2Ff83%2Ff8328b89-4f9f-4dec-baac-91
media%2F3c2%2F3c28aa10-0d36-4539-ac48-b0
media%2Fbd1%2Fbd1f37de-0e55-4664-a033-ff
media%2Ff9d%2Ff9d69b0d-3af5-478b-b1eb-9d
media%2Fa87%2Fa877555e-13de-48a8-9f25-c7
media%2Fdc8%2Fdc8d6fbd-cd36-4ff4-866c-0b
media%2F16e%2F16ef62b6-4f68-4d85-8cc8-c7
media%2F964%2F964b3d51-06c7-44ff-b838-ea
media%2Fbb0%2Fbb0503e7-6560-4c64-af6f-5a
Answer #1

1) a) f(x) = |x

b) g(x)=\left \{ g(x)\geq-2:\sqrt{x-2}-2 \right \}

c) h(x)=\frac{1}{x-1}-3

2) a) The square root function has been reflected horizontally about the line y=0 and has formula f(x)=-\sqrt{x} .

b) The square root function has been reflected vertically about the line x=0 and has formula f(x)=\sqrt{-x} .

3) a) Given, f(x) = -2x2

Now, f(-x) = -2(-x2) = -2x2 and -f(-x) = -[-2(-x)2] = 2x2

Since f(x) = f(-x), therefore the function is even.

b) Given f(x) = (1/x)+2x

Now, f(-x) = [1/(-x)]+2(-x) = -(1/x)-2x and -f(-x) = -[-(1/x)-2x] = (1/x)+2x

Since f(x) = -f(-x), therefore the function is odd.

c) Given f(x) = \sqrt{x}

Now, f(-x) = \sqrt{-x} and -f(-x) = -\sqrt{-x}

Since neither f(x) \neq f(-x) nor f(x) \neq -f(-x), therefore the function is neither even nor odd.

4) a) f(x) = -8x-x3

Now, f(-x) = -8(-x)-(-x)3 = 8x+x3 and -f(-x) = -(8x+x3) = -8x-x3

Since -f(-x) = f(x), therefore the function is odd.

b) f(x) = (x+1)2

Now, f(-x) = (-x+1)2 and -f(-x) = -(-x+1)2

Since neither f(x) \neq f(-x) nor f(x) \neq -f(-x), therefore the function is neither even nor odd.

c) f(x) = 7x6

Now, f(-x) = 7(-x)6 = 7x6 and -f(-x) = -7x6

Since f(x) = f(-x), therefore the function is even.

5) a) f(x)=3\left(x+1\right)^3-2

b) g(x)=\left|x\right|+2

6) a) f(x)=\left \{ f(x)\leq2:\sqrt{x-1}+2 \right \}

b) g(x)=|x+1|-1

7) a) f(x)=-2(x+2)^2

b) g(x)=\left \{ g(x)\leq0:\sqrt{8-4x} \right \}

8) a) f(x)=\left \{ f(x)\geq-2:\sqrt{4(x-1)}-2 \right \}

b) g(x)=-2\left(x+2\right)^3-1

9) a) g(x)=-\frac{1}{5}\sqrt{x}

b) g(x)=-3x^{1/3}

10) a) g(x)=\frac{1}{2x-4}+3

b) g(x)=6\sqrt{x-6}-9

11) a) g(x)=4\sqrt{x+5}+2

b) g(x)=\frac{1}{2\left(x+1\right)^2}-9

12) a) |x-11| = 3

b) |x-4| = 1/5

c) |x-10|\geq17

13) |x-80| = 0

14) |p-0.11|\leq0.025

15) |x-81|\leq15

16) |x-4.7|\leq0.012